Write a net ionic equation for the dissociation of hexacyanocobaltate (III) ion.

Write a net ionic equation for the reaction of cyanide ion with acid.

add the equations together to give the overall equation for the reaction that is described above.

the addition of acid to a solution of hexacyanocobaltate (III) ion produces hydrogen cyanide. this reaction can be expressed as the sum of 2 reactions: the dissociation of the complex ion and the reaction of cyanide ion with the acid.
Answer #1

Net ionic equation for the dissociation of hexacyanocobaltate (III) ion:

[Co(CN)6]-3 (aq) = Co+3 (aq) + 6 CN- (aq) ..........................Equation 1

Net ionic equation for the reaction of cyanide ion with acid :

CN -(aq) + H+ (aq) = HCN (aq)....................................Equation 2

Multiplying equation 2 by 6 and adding to equation 1; we get:

[Co(CN)6]-3 (aq) + 6 H+ (aq)= Co+3 (aq) + 6 HCN (aq) (This is the overall equation)